OpenStudy (anonymous):

If you have f(1) = 5 f(3) = 21 How can you find f'(3)?

jimthompson5910 (jim_thompson5910):

that's all you're given?

OpenStudy (anonymous):

Oh sorry i didnt see the other part on the next page O: For all real numbers of a and b, f(a+b) - f(a) = kab+2b^2

OpenStudy (anonymous):

A= 1 and B= 2

jimthompson5910 (jim_thompson5910):

ok figured that much, but what about k

jimthompson5910 (jim_thompson5910):

oh nvm, one sec

jimthompson5910 (jim_thompson5910):

f(a+b) - f(a) = kab+2b^2 f(1+2) - f(1) = k*1*2+2*2^2 f(3) - f(1) = k*1*2+2*2^2 21 - 5 = 2k - 8 solve for k to get k = ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (anonymous):

But you put minus 8 when it's positive O:

jimthompson5910 (jim_thompson5910):

oh my bad

jimthompson5910 (jim_thompson5910):

21 - 5 = 2k + 8 yeah k = 4

jimthompson5910 (jim_thompson5910):

so f(a+b) - f(a) = kab+2b^2 becomes f(a+b) - f(a) = 4ab+2b^2

jimthompson5910 (jim_thompson5910):

now if you factor out b from the right side, you will get this f(a+b) - f(a) = b(4a + 2b) then divide both sides by b to get [ f(a+b) - f(a) ]/b = 4a + 2b

jimthompson5910 (jim_thompson5910):

now if you take the limit of both sides as b ---> 0, then you will get the derivative f ' (a) replace a with 3 to get f ' (3)